OpenStudy (anonymous):

Find the sum of the infinite geometric series, if it exists. 4 - 1 +1/4 -1/16 + . . . A. - 1 B. 3 C. 16/5 D. does not exist

OpenStudy (anonymous):



OpenStudy (anonymous):

We have to find the common ratio since they told us this a geometric sequence. A common ratio (r) is basically the number divided by the preceding one. If we look at the sequence, we see that the common ration is -1/4 because: \[(-1) / (4) = -1/4\] \[(1/4)/(-1)= -1/4 \] \[(-1/16)/(1/4)= -1/4\] Now, there is a formula for finding the sum of an infinite geometric sequence. Basically, there is only one condition: \[\Large \left| r \right|<1\] Or in other words \[\Large -1<r<1\] r= -1/4 so it will work. The formula is \[\Large Sum~of~a~geometric~sequence~(S_\infty)=\] \[\Large \frac{ a_1 }{ 1-r }\] Where \[\Large a_1=first~term\] \[\Large r=~common~ratio\] Now plug in a_1= 4 and r= -1//4 in and tell me what you get

OpenStudy (anonymous):

Everywhere I said sequence, I meant series*

OpenStudy (courtneygraley009):

Use this math calculator to solve this geometric series. http://www.acalculator.com/quadratic-equation-calculator-formula-solver.html

OpenStudy (error1603):

b